# Fairness in Large Language Models: A Taxonomic Survey

### v1

@@ 1. Introduction
Large language models (LLMs) have demonstrated remarkable capabilities in addressing problems across diverse domains, ranging from chatbots (Glaese et al., 2022) to medical diagnoses (Wang et al., 2023g) and financial advisory (Shah et al., [n. d.]). Notably, their impact extends beyond fields directly associated with language processing, such as translation (Yao et al., 2023a) and text sentiment analysis (Nasukawa and Yi, 2003). LLMs also prove invaluable in broader applications including legal aid (Yu et al., 2022), healthcare (Singhal et al., 2023), and drug discovery (Sadybekov and Katritch, 2023). This highlights their adaptability and potential to streamline language-related tasks, making them indispensable tools across various industries and scenarios.
Despite their considerable achievements, LLMs may face fairness concerns stemming from biases inherited from the real-world and even exacerbate them (Zhao et al., 2019). Consequently, they could lead to discrimination against certain populations, especially in socially sensitive applications, across various dimensions such as race (An et al., 2022), age (Duan, [n. d.]), gender (Kotek et al., 2023), nationality (Venkit et al., 2023), occupation (Kirk et al., 2021), and religion (Abid et al., 2021). For instance, an investigation (Wan et al., 2023a) revealed that when tasked with generating a letter of recommendation for individuals named Kelly (a common female name) and Joseph (a common male name), ChatGPT, a prominent instance of LLMs, produced paragraphs describing Kelly and Joseph with random traits. Notably, Kelly was portrayed as warm and amiable (e.g., a well-regarded member), whereas Joseph was depicted as possessing greater leadership and initiative (e.g., a natural leader and role model). This observation indicates that LLMs tend to perpetuate gender stereotypes by associating higher levels of leadership with males.
@@ 2. An Overview Of The Taxonomy
As shown in Figure 1, we categorize recent studies on the fairness of LLMs according to three distinct perspectives: i) metrics for quantifying biases in LLMs, ii) algorithms for mitigating biases in LLMs, along with iii) resources for evaluating biases in LLMs. Regarding metrics for quantifying biases in LLMs, they are further categorized based on the data format used by metrics: i) embedding-based metrics, ii) probability-based metrics, and iii) generation-based metrics. Concerning bias mitigation techniques, they are structured according to the different stages within the LLMs workflow: i) pre-processing, ii) in-training, iii) intra-processing, and iv) post-processing. In addition, we collect resources for evaluating biases in LLMs and group them into Toolkits and Datasets. Specifically for Datasets, they are classified into two types based on the most appropriate metric type: i) probability-based and ii) generation-based.
@@ 3.1. Large Language Models
Language models are computational models with the capacity to comprehend and generate human language (Rosenfeld, 2000; Mikolov et al., 2010). The evolution of language models progresses from statistical language models to neural language models, pre-trained language models, and the current state of LLMs (Chu et al., 2024). Initial statistical language models, like N-gram models (Jelinek, 1998), estimate word likelihood based on the preceding context. However, N-gram models face challenges such as poor generalization ability, lack of long-term dependence, and difficulty capturing complex linguistic phenomena (Parsing, 2009). These limitations constrained the capabilities of language models until the emergence of transformers (Vaswani et al., 2017), which largely addressed these issues. Specifically, transformers became the backbone of modern language models (Wang et al., 2023f), attributable to their efficiency—an architecture free of recurrence that computes individual tokens in parallel—and effectiveness—attention facilitates spatial interaction across tokens dynamically dependent on the input itself. The advent of transformers has significantly expanded the scale of LLMs. These models not only demonstrate formidable linguistic capabilities but also rapidly approach human-level proficiency in diverse domains such as mathematics, reasoning, medicine, law, and programming (Bubeck et al., 2023). Nevertheless, LLMs frequently embed undesirable social stereotypes and biases, underscoring the emerging necessity to address such biases as a crucial undertaking.
@@ 3.2. Training Process of LLMs
Data preparation and preprocessing. The foundation of big language modeling is predicated on the availability of high-quality data. For LLMs, this entails the necessity of a vast corpus of textual data that is not only extensive but also rich in quality and diversity, which requires accurately represent the domain and language style that the model is aiming to grasp. Simultaneously, the datasets need to be large enough to provide sufficient training data for LLMs, and representative enough so that the models can adapt well to new and unseen texts (Sanh et al., 2021). Furthermore, the dataset needs to undergo a variety of processes, with data cleansing being a critical step involving the review and validation of data to eliminate discrimination and harmful content. For example, popular public sources for finding datasets, such as Kaggle [^1] https://www.kaggle.com/, Google Dataset Search [^2] https://datasetsearch.research.google.com/, Hugging Face [^3] https://huggingface.co/datasets, Data.gov [^4] https://data.gov/, and Wikipedia database [^5] https://en.wikipedia.org/wiki/Database, could all potentially harbor discriminatory content. This inclusion of biased information can adversely impact decision-making if fairness considerations are disregarded (Luo et al., 2023). Therefore, it is imperative to systematically remove any discriminatory content from the dataset to effectively reduce the risk of LLMs internalizing biased patterns.
Model selection and configuration. Most existing LLMs utilize transformer deep learning architectures, which have emerged as a preferred option for advanced natural language processing (NLP) tasks, such as Metas’s LLaMa (Touvron et al., 2023) and DeepAI’s GPT-3 (Brown et al., 2020). Several key elements of these models, such as the choice of loss function, the number of layers in transformer blocks, the number of attention heads, and various hyperparameters, need to be specified when configuring a transformer neural network. The configuration of these elements can vary depending on the desired use case and the characteristics of the training data. It is important to recognize that the model configuration directly influences the training duration and the potential introduction of bias during this process. One common source of bias amplification during the model training process is the selection of loss objectives mentioned above (Hovy and Prabhumoye, 2021). Typically, these objectives aim to enhance the accuracy of predictions. However, models may capitalize on chance correlations or statistical anomalies in the dataset to boost precision (e.g., all positive examples in the training data happened to come from male authors so that gender can be used as a discriminative feature) (Gururangan et al., 2018; Poliak et al., 2018). In essence, models may produce accurate results based on incorrect rationales, resulting in discrimination.
@@
Alignment with human. During training, the model is exposed to examples such as “what is the capital of India?” paired with the labeled output “Delhi”, enabling it to learn the relationship between input queries and expected output responses. This equips the model to accurately answer similar questions, like “What is the capital of France?” resulting in the answer “Paris”. While this highlights the model’s capabilities, there are scenarios where its performance may falter, particularly when queried like “Whether men or women are better leaders?” where the model may generate biased content. This introduces concerns about bias in the model’s responses. For this purpose, InstructGPT (Ouyang et al., 2022) designs an effective tuning approach that enables LLMs to follow the expected instructions, which utilizes the technique of reinforcement learning with human feedback (RLHF) (Christiano et al., 2017; Ouyang et al., 2022). RLHF is a ML technique that uses human feedback to optimize LLMs to self-learn more efficiently. Reinforcement learning techniques train model to make decisions that maximize rewards, making their outcomes more accurate. RLHF incorporates human feedback in the rewards function, so the LLMs can perform tasks more aligned with human values such as helpfulness, honesty, and harmlessness. Notably, ChatGPT is developed based on a similar technique as InstructGPT, exhibits a strong ability to generate high-quality, benign responses, including the ability to avoid engaging with offensive queries.
@@ 3.3. Factors Contributing to Bias in LLMs
iii) Label bias. In instruction tuning scenarios, biases can arise from the subjective judgments of human annotators who provide labels or annotations for training data (Sap et al., 2019). This occurs when annotators inject their personal beliefs, perspectives, or stereotypes into the labeling process, inadvertently introducing bias into the model. Another potential source of bias is the RLHF approach discussed in Section 3, where human feedback is used to align LLMs with human values. While this method aims to improve model behavior by incorporating human input, it inevitably introduces subjective notions into the feedback provided by human. These subjective ideas can influence the model’s training and decision-making processes, potentially leading to biased outcomes. Therefore, it is crucial to implement measures to detect and mitigate bias when performing instruction tuning, such as diversifying annotator perspectives, and evaluating model performance using fairness metrics.
@@ 4.1. Group Fairness
Existing fairness definitions (Hardt et al., 2016; Dwork et al., 2012) at the group level aim to emphasize that algorithmic decisions neither favor nor harm certain subgroups defined by the sensitive attribute, which often derives from legal standards or topics of social sensitivity, such as gender, race, religion, age, sexuality, nationality, and health conditions. These attributes delineate a variety of demographic or social groups, with sensitive attribute categorized as either binary (e.g., male, female) or pluralistic (e.g., Jewish, Islamic, Christian). However, existing fairness metrics, developed primarily for traditional machine learning tasks (e.g., classification), rely on the availability of clear class labels and corresponding numbers of members belonging to each demographic group for quantification. For example, when utilizing the German Credit Dataset (Asuncion and Newman, 2007) and considering the relationship between gender and credit within the framework of statistical parity (where the probability of granting a benefit, such as credit card approval, is the same for different demographic groups) (Verma and Rubin, 2018), machine learning algorithms like decision trees can directly produce a binary credit score for each individual. This enables the evaluation of whether there is an equal probability for male and female applicants to obtain a good predicted credit score. However, this quantification presupposes the applicability of class labels and relies on the number of members from different demographic groups belonging to each class label, an assumption that does not hold for LLMs. LLMs, which are often tasked with generative or interpretive functions rather than simple classification, necessitate a different linguistic approach to such demographic group-based disparities; Instead of direct label comparison, group fairness in LLMs involves ensuring that word embeddings, vector representations of words or phrases, do not encode biased associations. For example, the embedding for “doctor” should not be closer to male-associated words than to female-associated ones. This would indicate that the LLM associates both genders equally with the profession, without embedding any societal biases that might suggest one gender is more suited to the profession than the other.
@@ 5.1. Embedding-based Metrics
(1) $s(\boldsymbol{w},A,B)=\frac{1}{n}\sum_{\boldsymbol{a}\in A}\cos(\boldsymbol{w},\boldsymbol{a})-\frac{1}{n}\sum_{\boldsymbol{b}\in B}\cos(\boldsymbol{w},\boldsymbol{b})\vspace{-0.1cm}$
Subsequently, to quantify bias in the sets $X$ and $Y$, the effect size is used as a normalized measure for the association difference between the target sets:
(2) $WEAT(X,Y,A,B)=\frac{\text{ mean }_{\boldsymbol{x}\in X}s(\boldsymbol{x},A,B)-\text{ mean }_{\boldsymbol{y}\in Y}s(\boldsymbol{y},A,B)}{\operatorname{stddev}_{\boldsymbol{w}\in X\cup Y}s(\boldsymbol{w},A,B)}$
where $\text{mean}_{\boldsymbol{x}\in X}s(\boldsymbol{x},A,B)$ represents the average of $s(x,A,B)$for $x$ in $X$, while $\operatorname{stddev}_{\boldsymbol{w}\in X\cup Y}s(\boldsymbol{w},A,B)$ denotes the standard deviation across all word biases of $x$ in $X$.
@@ 5.2. Probability-based Metrics
Probability-based metrics formalize bias by analyzing the probabilities assigned by LLMs to various options, often predicting words or sentences based on templates (Bartl et al., 2020; Rudinger et al., 2018) or evaluation sets (Nangia et al., 2020; Felkner et al., 2023). These metrics are generally divided into two categories: masked tokens, which assess token probabilities in fill-in-the-blank templates, and pseudo-log-likelihood is utilized to assess the variance in probabilities between counterfactual pairs of sentences.
@@
Log Probability Bias Score (LPBS) (Kurita et al., 2019). LPBS adopts template sentences similar to DisCO. However, unlike DisCO, LPBS corrects for the influence of inconsistent prior probabilities of target attributes. Specifically, for computing the association between the target gender male and the attribute doctor, LPBS first feeds the masked sentence “[MASK] is a doctor” into the model to obtain the probability of the sentence “he is a doctor”, denoted as $P_{tar_{male}}$. Then, to correct for the influence of inconsistent prior probabilities of target attributes, LPBS feeds the masked sentence “[MASK] is a [MASK]” into the model to obtain the probability of the sentence “he is a [MASK]”, denoted as $P_{pri_{male}}$. This process is repeated with “he” replaced by “she” for the target gender female. Finally, the bias is assessed by comparing the normalized probability scores for two contrasting attribute words and the specific formula is defined as:
@@ 5.3. Generation-based Metrics
(5) $\operatorname{bias}(w)=\log(\frac{P(w\mid female)}{P(w\mid male)}),P(w\mid g)=\frac{d(w,g)/\Sigma_{i}d\left(w_{i},g\right)}{d(g)/\Sigma_{i}d\left(w_{i}\right)}$
@@ 6.1. Pre-processing
ii) Prompt Tuning. In contrast to CDA, prompt tuning (Lester et al., 2021) focuses on reducing biases in LLMs by refining prompts provided by users. Prompt tuning can be categorized into two types: hard prompts and soft prompts. The former refers to predefined prompts that are static and may be considered as templates. Although templates provide some flexibility, the prompt itself remains mostly unchanged, hence the term “hard prompt.” On the other hand, soft prompts are created dynamically during the prompt tuning process. Unlike hard prompts, soft prompts cannot be directly accessed or edited as text. Soft prompts are essentially embeddings, a series of numbers, that contain information extracted from the broader model. As a specific example of hard prompt, Mattern et al. (Mattern et al., 2022) introduced an approach focusing on analyzing the bias mitigation effects of prompts across various levels of abstraction. In their experiments, they observed that the effects of debiasing became more noticeable as prompts became less abstract, as these prompts encouraged GPT-3 to utilize gender-neutral pronouns more frequently. In terms of soft prompt method, Fatemi et al. (Fatemi et al., 2021) focus on achieving gender equality by freezing model parameters and utilizing gender-neutral datasets to update biased word embeddings associated with occupations, effectively reducing bias in prompts. Overall, the disadvantage of hard prompts is their lack of flexibility, while the drawback of soft prompts is the lack of interpretability.

### v2

@@ 1. Introduction
Large language models (LLMs) have demonstrated remarkable capabilities in addressing problems across diverse domains, ranging from chatbots (Glaese et al., 2022) to medical diagnoses (Wang et al., 2023h) and financial advisory (Shah et al., [n. d.]). Notably, their impact extends beyond fields directly associated with language processing, such as translation (Yao et al., 2023a) and text sentiment analysis (Nasukawa and Yi, 2003). LLMs also prove invaluable in broader applications including legal aid (Yu et al., 2022), healthcare (Singhal et al., 2023), and drug discovery (Sadybekov and Katritch, 2023). This highlights their adaptability and potential to streamline language-related tasks, making them indispensable tools across various industries and scenarios.
Despite their considerable achievements, LLMs may face fairness concerns stemming from biases inherited from the real world and even exacerbate them (Zhao et al., 2019). Consequently, they could lead to discrimination against certain populations, especially in socially sensitive applications, across various dimensions such as race (An et al., 2022), age (Duan, [n. d.]), gender (Kotek et al., 2023), nationality (Venkit et al., 2023), occupation (Kirk et al., 2021), and religion (Abid et al., 2021). For instance, an investigation (Wan et al., 2023a) revealed that when tasked with generating a letter of recommendation for individuals named Kelly (e.g., a common female name) and Joseph (e.g., a common male name), ChatGPT, a prominent instance of LLMs, produced paragraphs describing Kelly and Joseph with random traits. Notably, Kelly was portrayed as warm and amiable (e.g., a well-regarded member), whereas Joseph was depicted as possessing greater leadership and initiative (e.g., a natural leader and role model). This observation indicates that LLMs tend to perpetuate gender stereotypes by associating higher levels of leadership with males.
@@ 2. An Overview Of The Taxonomy
As shown in Figure 1, we categorize recent studies on the fairness of LLMs according to three distinct perspectives: i) metrics for quantifying biases in LLMs, ii) algorithms for mitigating biases in LLMs, and iii) resources for evaluating biases in LLMs. Regarding metrics for quantifying biases in LLMs, they are further categorized based on the data format used by metrics: i) embedding-based metrics, ii) probability-based metrics, and iii) generation-based metrics. Concerning bias mitigation techniques, they are structured according to the different stages within the LLMs workflow: i) pre-processing, ii) in-training, iii) intra-processing, and iv) post-processing. In addition, we collect resources for evaluating biases in LLMs and group them into Toolkits and Datasets. Specifically for Datasets, they are classified into two types based on the most appropriate metric type: i) probability-based and ii) generation-based.
@@ 3.1. Large Language Models
Language models are computational models with the capacity to comprehend and generate human language (Rosenfeld, 2000; Mikolov et al., 2010). The evolution of language models progresses from statistical language models to neural language models, pre-trained language models, and the current state of LLMs (Chu et al., 2024a). Initial statistical language models, like N-gram models (Jelinek, 1998), estimate word likelihood based on the preceding context. However, N-gram models face challenges such as poor generalization ability, lack of long-term dependence, and difficulty capturing complex linguistic phenomena (Parsing, 2009). These limitations constrained the capabilities of language models until the emergence of transformers (Vaswani et al., 2017), which largely addressed these issues. Specifically, transformers became the backbone of modern language models (Wang et al., 2023g), attributable to their efficiency—an architecture free of recurrence that computes individual tokens in parallel—and effectiveness—attention facilitates spatial interaction across tokens dynamically dependent on the input itself. The advent of transformers has significantly expanded the scale of LLMs. These models not only demonstrate formidable linguistic capabilities but also rapidly approach human-level proficiency in diverse domains such as mathematics, reasoning, medicine, law, and programming (Bubeck et al., 2023). Nevertheless, LLMs frequently embed undesirable social stereotypes and biases, underscoring the emerging necessity to address such biases as a crucial undertaking.
@@ 3.2. Training Process of LLMs
Data preparation and preprocessing. The foundation of big language modeling is predicated on the availability of high-quality data. For LLMs, this entails the necessity of a vast corpus of textual data that is not only extensive but also rich in quality and diversity, which requires accurately representing the domain and language style that the model is aiming to grasp. Simultaneously, the datasets need to be large enough to provide sufficient training data for LLMs, and representative enough so that the models can adapt well to new and unseen texts (Sanh et al., 2021). Furthermore, the dataset needs to undergo a variety of processes, with data cleansing being a critical step involving the review and validation of data to eliminate discrimination and harmful content. For example, popular public sources for finding datasets, such as Kaggle [^1] https://www.kaggle.com/, Google Dataset Search [^2] https://datasetsearch.research.google.com/, Hugging Face [^3] https://huggingface.co/datasets, Data.gov [^4] https://data.gov/, and Wikipedia database [^5] https://en.wikipedia.org/wiki/Database, could all potentially harbor discriminatory content. This inclusion of biased information can adversely impact decision-making if fairness considerations are disregarded (Luo et al., 2023). Therefore, it is imperative to systematically remove any discriminatory content from the dataset to effectively reduce the risk of LLMs internalizing biased patterns.
Model selection and configuration. Most existing LLMs utilize transformer deep learning architectures, which have emerged as a preferred option for advanced natural language processing (NLP) tasks, such as Metas’s LLaMa (Touvron et al., 2023) and DeepAI’s GPT-3 (Brown et al., 2020). Several key elements of these models, such as the choice of the loss function, the number of layers in transformer blocks, the number of attention heads, and various hyperparameters, need to be specified when configuring a transformer neural network. The configuration of these elements can vary depending on the desired use case and the characteristics of the training data. It is important to recognize that the model configuration directly influences the training duration and the potential introduction of bias during this process. One common source of bias amplification during the model training process is the selection of loss objectives mentioned above (Hovy and Prabhumoye, 2021). Typically, these objectives aim to enhance the accuracy of predictions. However, models may capitalize on chance correlations or statistical anomalies in the dataset to boost precision (e.g., all positive examples in the training data happened to come from male authors so that gender can be used as a discriminative feature) (Gururangan et al., 2018; Poliak et al., 2018). In essence, models may produce accurate results based on incorrect rationales, resulting in discrimination.
@@
Alignment with human. During training, the model is exposed to examples such as “What is the capital of India?” paired with the labeled output “Delhi,” enabling it to learn the relationship between input queries and expected output responses. This equips the model to accurately answer similar questions, like “What is the capital of France?” resulting in the answer “Paris”. While this highlights the model’s capabilities, there are scenarios where its performance may falter, particularly when queried like “Whether men or women are better leaders?” where the model may generate biased content. This introduces concerns about bias in the model’s responses. For this purpose, InstructGPT (Ouyang et al., 2022) designs an effective tuning approach that enables LLMs to follow the expected instructions, which utilizes the technique of reinforcement learning with human feedback (RLHF) (Christiano et al., 2017; Ouyang et al., 2022). RLHF is an ML technique that uses human feedback to optimize LLMs to self-learn more efficiently. Reinforcement learning techniques train the model to make decisions that maximize rewards, making their outcomes more accurate. RLHF incorporates human feedback in the rewards function, so the LLMs can perform tasks more aligned with human values such as helpfulness, honesty, and harmlessness. Notably, ChatGPT is developed based on a similar technique as InstructGPT and exhibits a strong ability to generate high-quality, benign responses, including the ability to avoid engaging with offensive queries.
@@ 3.3. Factors Contributing to Bias in LLMs
iii) Label bias. In instruction tuning scenarios, biases can arise from the subjective judgments of human annotators who provide labels or annotations for training data (Sap et al., 2019). This occurs when annotators inject their personal beliefs, perspectives, or stereotypes into the labeling process, inadvertently introducing bias into the model. Another potential source of bias is the RLHF approach discussed in Section 3, where human feedback is used to align LLMs with human values. While this method aims to improve model behavior by incorporating human input, it inevitably introduces subjective notions into the feedback provided by humans. These subjective ideas can influence the model’s training and decision-making processes, potentially leading to biased outcomes. Therefore, it is crucial to implement measures to detect and mitigate bias when performing instruction tuning, such as diversifying annotator perspectives, and evaluating model performance using fairness metrics.
@@ 4.1. Group Fairness
Existing fairness definitions (Hardt et al., 2016; Dwork et al., 2012) at the group level aim to emphasize that algorithmic decisions neither favor nor harm certain subgroups defined by the sensitive attribute, which often derives from legal standards or topics of social sensitivity, such as gender, race, religion, age, sexuality, nationality, and health conditions. These attributes delineate a variety of demographic or social groups, with sensitive attributes categorized as either binary (e.g., male, female) or pluralistic (e.g., Jewish, Islamic, Christian). However, existing fairness metrics, developed primarily for traditional machine learning tasks (e.g., classification), rely on the availability of clear class labels and corresponding numbers of members belonging to each demographic group for quantification. For example, when utilizing the German Credit Dataset (Asuncion and Newman, 2007) and considering the relationship between gender and credit within the framework of statistical parity (where the probability of granting a benefit, such as credit card approval, is the same for different demographic groups) (Verma and Rubin, 2018), machine learning algorithms like decision trees can directly produce a binary credit score for each individual. This enables the evaluation of whether there is an equal probability for male and female applicants to obtain a good predicted credit score. However, this quantification presupposes the applicability of class labels and relies on the number of members from different demographic groups belonging to each class label, an assumption that does not hold for LLMs. LLMs, which are often tasked with generative or interpretive functions rather than simple classification, necessitate a different linguistic approach to such demographic group-based disparities; Instead of direct label comparison, group fairness in LLMs involves ensuring that word embeddings, vector representations of words or phrases, do not encode biased associations. For example, the embedding for “doctor” should not be closer to male-associated words than to female-associated ones. This would indicate that the LLM associates both genders equally with the profession, without embedding any societal biases that might suggest one gender is more suited to the profession than the other.
@@ 5.1. Embedding-based Metrics
(1) $s(\boldsymbol{w},A,B)=\frac{1}{n}\sum_{\boldsymbol{a}\in A}\cos(\boldsymbol{w},\boldsymbol{a})-\frac{1}{n}\sum_{\boldsymbol{b}\in B}\cos(\boldsymbol{w},\boldsymbol{b})$
Subsequently, to quantify bias in the sets $M$ and $F$, the effect size is used as a normalized measure for the association difference between the target sets:
(2) $WEAT(M,F,A,B)=\frac{\text{ mean }_{\boldsymbol{m}\in M}s(\boldsymbol{m},A,B)}{\operatorname{stddev}_{\boldsymbol{w}\in M\cup F}s(\boldsymbol{w},A,B)}$
[…]
where $\text{mean}_{\boldsymbol{m}\in M}s(\boldsymbol{m},A,B)$ represents the average of $s(m,A,B)$for $m$ in $M$, while $\operatorname{stddev}_{\boldsymbol{w}\in M\cup F}s(\boldsymbol{w},A,B)$ denotes the standard deviation across all word biases of $m$ in $M$.
@@ 5.2. Probability-based Metrics
Probability-based metrics formalize bias by analyzing the probabilities assigned by LLMs to various options, often predicting words or sentences based on templates (Bartl et al., 2020; Rudinger et al., 2018) or evaluation sets (Felkner et al., 2023). These metrics are generally divided into two categories: masked tokens, which assess token probabilities in fill-in-the-blank templates, and pseudo-log-likelihood is utilized to assess the variance in probabilities between counterfactual pairs of sentences.
@@
Log Probability Bias Score (LPBS) (Kurita et al., 2019). LPBS adopts template sentences similar to DisCO. However, unlike DisCO, LPBS corrects for the influence of inconsistent prior probabilities of target attributes. Specifically, for computing the association between the target gender male and the attribute doctor, LPBS first feeds the masked sentence “[MASK] is a doctor” into the model to obtain the probability of the sentence “he is a doctor”, denoted as $P_{tar_{male}}$. Then, to correct for the influence of inconsistent prior probabilities of target attributes, LPBS feeds the masked sentence “[MASK] is a [MASK]” into the model to obtain the probability of the sentence “he is a [MASK]”, denoted as $P_{pri_{male}}$. This process is repeated with “he” replaced by “she” for the target gender female. Finally, the bias is assessed by comparing the normalized probability scores for two contrasting attribute words, and the specific formula is defined as:
@@ 5.3. Generation-based Metrics
(5) $\operatorname{bias}(w)=\log(\frac{P(w\mid female)}{P(w\mid male)}),P(w\mid g)=\frac{d(w,g)/\Sigma_{i}d\left(w_{i},g\right)}{d(g)/\Sigma_{i}d\left(w_{i}\right)}\vspace{-0.1cm}$
@@ 6.1. Pre-processing
ii) Prompt Tuning. In contrast to CDA, prompt tuning (Lester et al., 2021) focuses on reducing biases in LLMs by refining prompts provided by users. Prompt tuning can be categorized into two types: hard prompts and soft prompts. The former refers to predefined prompts that are static and may be considered as templates. Although templates provide some flexibility, the prompt itself remains mostly unchanged, hence the term “hard prompt.” On the other hand, soft prompts are created dynamically during the prompt tuning process. Unlike hard prompts, soft prompts cannot be directly accessed or edited as text. Soft prompts are essentially embeddings, a series of numbers, that contain information extracted from the broader model. As a specific example of a hard prompt, Mattern et al. (Mattern et al., 2022) introduced an approach focusing on analyzing the bias mitigation effects of prompts across various levels of abstraction. In their experiments, they observed that the effects of debiasing became more noticeable as prompts became less abstract, as these prompts encouraged GPT-3 to utilize gender-neutral pronouns more frequently. In terms of soft prompt method, Fatemi et al. (Fatemi et al., 2021) focus on achieving gender equality by freezing model parameters and utilizing gender-neutral datasets to update biased word embeddings associated with occupations, effectively reducing bias in prompts. Overall, the disadvantage of hard prompts is their lack of flexibility, while the drawback of soft prompts is the lack of interpretability.